OpenStudy (anonymous):

On the top of the Sears Building in Chicago, is a TV tower. From a point 400 ft. away the bottom of the tower is at an angle of elevation equal to 38 degrees and the top of the tower is an angle of elevation equal to 42 degrees. How tall is the tower?

OpenStudy (anonymous):

tan 38 =x/400, tan 42 = y/400, the height would be y - x